Circular No. 6758

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

CATACLYSMIC VARIABLE IN AURIGA Y.-l. Qiu, Q.-y. Qiao, W.-d. Li, and J.-y. Hu, Beijing Astronomical Ob-

servatory (BAO); and A. Esamdin, Urumqi Astronomical Station, report their discovery, in the course of BAO supernova survey, of a new cataclysmic variable star at α = 7^{h}27^{m}52^{s}.20, δ = +40^{o}46′55′′.5 (equinox 2000.0). Estimated R magnitudes obtained with 0.60-m telescope at Xinglong station: Oct. 10 UT, [18.5; 19.82, 15.7; 20.77, 15.9. The original Palomar Sky Survey plate shows no object at the position of this star. A spectrogram of the variable was obtained with the 2.16-m telescope at Xinglong on Oct. 20.85. The spectrum shows a very blue continuum with absorption lines of Hδ, Hγ, Hβ, and He I (468.6 nm). Hβ and He I (468.6 nm) display weak emission in the absorption core. The spectrum is similar to that of the dwarf nova CY Lyr at maximum.

NOVA IN THE LARGE MAGELLANIC CLOUD 1997 W. Liller, Viña del Mar, Chile, reports that his films taken on June

5.97 and 25.96 UT (limiting magnitude ∼ 13.5) show no evidence of the nova reported by the MACHO collaboration on IAUC 6756.

COMET C/1997 T1 (UTSUNOMIYA) Extension to ephemeris on IAUC 6753 (elements MPC 30738):

1997TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 19 19^{h}45.^{m}59 +54^{o}39.^{′}7 1.043 1.557 99.^{o}6 39.^{o}1 10.0
24 19 23.66 +47 12.3 1.095 1.524 93.5 40.7 10.0
29 19 10.12 +40 26.7 1.166 1.493 87.1 41.7 10.1
Nov. 3 19 01.45 +34 33.0 1.250 1.464 80.7 42.0 10.1
8 18 55.77 +29 29.9 1.342 1.439 74.5 41.6 10.2
13 18 52.02 +25 12.0 1.439 1.417 68.5 40.5 10.3
18 18 49.57 +21 32.3 1.537 1.398 62.9 39.0 10.4
23 18 48.03 +18 24.3 1.634 1.383 57.5 37.0 10.5
28 18 47.14 +15 42.4 1.729 1.371 52.4 34.8 10.6
Dec. 3 18 46.71 +13 21.8 1.819 1.364 47.6 32.2 10.6
8 18 46.62 +11 18.8 1.903 1.360 43.0 29.6 10.7
13 18 46.77 + 9 30.3 1.980 1.361 38.8 27.0 10.8
18 18 47.08 + 7 53.8 2.049 1.365 35.0 24.4 10.9
1997 October 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green